An angle measures six degrees more than its complement. Find the angle measures.
Angle = _______
Complement = _______
x = ______

Answer by waynest(281)

x + (x + 6) = 90
2x + 6 = 90
2x + 6 - 6 = 90 - 6
2x = 84
2x/2 = 84/2
x = 42
angle: 42 + 6 = 48
complement: 42
x = 42